OpenStudy (anonymous):

5. Solve for k. k + (–45) = –18 A. –27 B. –63 C. 27 D. 63

OpenStudy (anonymous):

k+(-45)=-18 @iGreen help me please

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

HELP

OpenStudy (igreen):

k + (–45) = –18 k - 45 = -18 Add 45 to both sides: What's 45 + -18? @lloio

OpenStudy (texaschic101):

k + (-45) = -18 --- a positive times a neg = a neg k - 45 = -18 -- add 45 to both sides k = -18 + 45 k = 27

OpenStudy (igreen):

@texaschic101 I got warned for giving out direct answers..I told them that I also gave an explanation..but he says you're supposed to lead them and NOT to give them the answer..even though their Code of Conduct says to not JUST give them the answer..it's stupid.

OpenStudy (anonymous):

+27

OpenStudy (texaschic101):

I know...I shouldn't do that. I have been banned before for doing it. I sometimes just help people out. Thats probably why I do not even try to become a mod.

OpenStudy (anonymous):

\[45+ -18=27\]

OpenStudy (igreen):

Lol, Ik

OpenStudy (anonymous):

MIND HELPING ME WITH MORE? sorry about cap locks

OpenStudy (texaschic101):

what ya got

OpenStudy (texaschic101):

oh crap...I lost track of time. I have to go now, but I will be back on later...sorry

OpenStudy (anonymous):

;Paul biked 14.6 miles on Saturday and 12.8 miles on Sunday. Mariah biked the same total distance as Paul, but only biked 11.7 miles on Saturday. How far did Mariah bike on Sunday? A. 15.7 miles B. 27.8 miles C. 13.5 miles D. 1.8 miles

OpenStudy (anonymous):

@iGreen help

OpenStudy (texaschic101):

Paul biked (14.6 + 12.8) = 27.4 Mariah biked same total distance (27.4)....but biked 11.7 on Sat. So she biked (27.4 - 11.7) = 15.7 miles on Sunday I got to go now...c ya later